OpenStudy (anonymous):

for a positive integer x, 10 percent of x percent of 1000 is equal to what? teh possible anwsers are "X" "10x" "100x" "1000x" plzz help:s

OpenStudy (anonymous):

lets make \(x=7\) then 10% of 7% of 1,000 is \(0.1\times 0.07\times 1000=7\)

OpenStudy (anonymous):

thank you for your help;d i didn't think of it that way.. and i'm gessing it works on all numbers right?

OpenStudy (anonymous):

yes since \(10\%\) of \(x\%\) of \(1000\) is \[\frac{1}{10}\times \frac{x}{100}\times 1000=\frac{x}{1000}\times 1000=x\]

OpenStudy (anonymous):

okkkk thank you very much:D